OpenStudy (anonymous):

Points A and B lie on the graph of the equation y = 7x. What are the y-coordinates for each point? Point A: (−2, ) Point B: (4, )

OpenStudy (coconutjj):

sub x coord into 7x

OpenStudy (ddcamp):

For point A, x = -2 (the left number in the ordered pair) Since y = 7x, for point A, y = 7 * x = 7 * (-2) = -14 Can you find y for point B?

OpenStudy (anonymous):

@DDCamp 28?